OpenStudy (anonymous):

completing the square x^2-8x+32=o

OpenStudy (anonymous):

i think i just did this one was it wrong?

OpenStudy (anonymous):

x^2-2*4x+(2)^2+32=4^2 (x-4)^2=16-32 (x-4)^2=-16 x-4=+-4i x=4(1+-i)